OpenStudy (anonymous):

How do you find the inverse of f(x)= 12x-3 over 4x-8

OpenStudy (anonymous):

y = (12x-3)/(4x-8) (4x-8)y = 12x - 3 4xy - 8y = 12x - 3 4xy - 12x = 8y - 3 4x(y-3) = 8y - 3 4x = (8y - 3)/(y-3) x = (1/4)[ (8y-3)/(y-3) ] Now switch x and y to find the inverse... f inverse = (1/4)[ (8x-3)/(x-3) ]

OpenStudy (anonymous):

thanks